OpenStudy (anonymous):

mc038-1.jpg What is the value of x to the nearest tenth?

OpenStudy (anonymous):

@kidrah69

OpenStudy (wolf1728):

ally - you are familiar with the Pythagorean Theorem?

OpenStudy (anonymous):

yes a^2+b^2=c^2

OpenStudy (wolf1728):

You can apply that to this problem to get the value of x

OpenStudy (anonymous):

would I do, x^2+4.7^2=18.8^2

OpenStudy (wolf1728):

You have to determine side1 side 2 and the hypotenuse. x would be the hypotenuse 4.7 would be one side what would be the other side?

OpenStudy (anonymous):

18.8

OpenStudy (wolf1728):

Look carefully. Is 18.8 a "side" of the triangle?

OpenStudy (anonymous):

ohhhhhh 9.4

OpenStudy (wolf1728):

Very good !

OpenStudy (wolf1728):

So x^2 = 9.4^2 + 4.7^2

OpenStudy (anonymous):

10.5?

OpenStudy (wolf1728):

Yes - or to be really accurate 10.5095194942

OpenStudy (anonymous):

Thank you!

OpenStudy (wolf1728):

u r welcome :-)